Suppose that the following equations govern planned spending in the US:
C = 500 + 0.75(Y-T)
T = 0.2Y – 800
I = 3000 – 64000r
G = 3200
NX = 1000 – 10e (e =“trade weighted” real ex. rate. As always, increase in e = $ appreciation)
NFO = 500 – 60000(r – r
FOR)
r
FOR = 3%
a) Explain how NFO responds to an increase in the Home interest rate, and an increase
r
FOR, based on the equation. What economics story does this coefficient represent?
b) We’re going to look at an increase in the home interest rate from 2.5% to 3%. First,
let’s take a look at the new international piece of the model. Calculate NFO, NX,
and the exchange rate for each value of the interest rate. Based on these numbers,
draw the NFO=NX graph for the interest rate increase, and talk through the
economics: how the change in interest rates changes NFO, and how that leads to
changes in the exchange rate and NX.
c) Now, combine the equations above to find an expression for AE, then impose the
Y=AE expression to derive the IS curve. As in lecture, work with AE = C + I + G
+NFO to end up with Y=f(r). Use your IS curve to calculate the level of short-run
GDP for the original interest rate of 2.5%, and the new interest rate of 3%.
d) Use the multiplier math framework (Final Chg GDP = Multiplier x Initial Chg in
Spending) to explain this change in GDP? The change in NX from part b is part of
this story, but don’t forget about our pre-PS5 domestic interest rate story – where
else does r show up?
e) It is often said that “monetary policy is stronger in an open economy.” Explain this
statement based on the “domestic” and “international” changes in GDP from part
d). I don’t just want the numbers side of things – what is the underlying economics
story for why monetary policy is more effective when the economy has an
international sector?
Answer #1

NFO 500-60,000 CFL-3) NFO = Sbo-600 NX = 1900-10e 1仁D 500-60000亿ー 1000-0 e ウ 7 }Q. D500-1000-60000九华一10e 6000 tire is an appe2 P2 NPO, NX hy 2.ham 0.4Y = 18 4800-60000ん = 462000-150000ん ST y 62000-Y58000 shat yumthere is ニー60000 △ 4足 Cz 0.75 0,75(-0.2)a) NFO = 500 – 60000(r – rFOR)

from the above equation it can be seen

1.there is a negative relation between NFO and r

2. there is positive relation between NFO and rFOR

1 means if there is an increase in the home interest rate , NFO will decrease . This is because if the home or domestic interest rates are higher, the funds or capital would flow from foreign countries to the home country .So, the outflows will decrease.

2 on the contrary means if there is an decrease in the foreign interest rate, NFO will increase . This is because then high foreign interest rate will attract capital to foreign countries. So, the outflows will increase.

The absolute coefficient ( 60000) represents the sensitivity of NFO to change in r or rFOR.

It states for every 1 percent point increase in r (home interest rate) , NFO would be lower by 60000. Similarly, For or every 1 percent point increase in rFOR (foreign interest rate) , NFO would be higher by 60000.